C6 H12 O6 + 2C2 H5 OH+ 2CO2
if 3.44 mol of ethanol are produced during the reaction, how many glucose have been used?
1
Expert's answer
2020-10-10T09:39:40-0400

Solution:

The chemical equation for alcoholic fermentation is:

C6H12O6 → 2C2H5OH + 2CO2

According to the equation: n(C6H12O6) = n(C2H5OH)/2

Hence,

n(C6H12O6) = 3.44 mol / 2 = 1.72 mol


The molar mass of glucose (C6H12O6) is 180.156 g/mol.

m(C6H12O6) = n(C6H12O6) × M(C6H12O6) = 1.72 mol × 180.156 g/mol = 309.87 g


1.72 mol of glucose corresponds to 309.87 g of glucose.


Answer: 1.72 mol of glucose have been used that corresponds to 309.87 g of glucose.